OpenStudy (thunder33):

What is the probability of getting heads when flipping a coin and getting a number greater than or equal to 4 when rolling a single die?

OpenStudy (adonine):

Assuming that the coin is 'fair' (i.e. a regular coin), the classical interpretation is that the probability of getting heads is exactly 0.5. Again, assuming fair die, the probability of getting any value when rolling is uniformly distributed, so 1/6 for any given value. Since there are two values that are greater than or equal to 5 on a fair die, the probability of getting a value greater than or equal to 5 is 1/6 + 1/6 = 2/6 = 1/3. The probability of both happening, is the product of both probabilities: P = (1/2)*(1/3) = 1/6